How much faster does He atom effuse than N₂ molecule at the same temperature? (MMHe = 2g, MMN₂ = 28g)


1
Expert's answer
2022-05-24T06:08:03-0400

The rate of effusion of a gas is inversely proportional to the square root of the molecular weight of the gas. The lighter a gas is, the faster it will effuse; the heavier a gas is, the slower it will effuse:

"\frac{Effuse–rate_1} {Effuse–rate_2} =\sqrt{\frac {M_2} {M_1}}=\sqrt{ \frac {28}{2}}=\sqrt{14}=3,74"

ANSWER: The effusion of the He atom is 3.74 times faster than the effusion of the N2 molecule.